### v1

@@ Abstract
Can a mere next-token predictor faithfully model human intelligence? We crystallize this intuitive concern, which is fragmented in the literature. As a starting point, we argue that the two often-conflated phases of next-token prediction — autoregressive inference and teacher-forced training — must be treated distinctly. The popular criticism that errors can compound during autoregressive inference, crucially assumes that teacher-forcing has learned an accurate next-token predictor. This assumption sidesteps a more deep-rooted problem we expose: in certain classes of tasks, teacher-forcing can simply fail to learn an accurate next-token predictor in the first place. We describe a general mechanism of how teacher-forcing can fail, and design a minimal planning task where both the Transformer and the Mamba architecture empirically fail in that manner — remarkably, despite the task being straightforward to learn. We provide preliminary evidence that this failure can be resolved when training to predict multiple tokens in advance. We hope this finding can ground future debates and inspire explorations beyond the next-token prediction paradigm. We make our code available under https://github.com/gregorbachmann/Next-Token-Failures
@@ 1 Introduction
Long after its inception in the seminal work of Shannon (1948; 1951), next-token prediction has made its way into becoming a core part of the modern language model. But despite its long list of achievements, there is a small but growing belief that a next-token predicting model is merely an impressive improv artist that cannot truly model human thought. Humans, when navigating the world, meticulously imagine, curate and backtrack plans in their heads before executing them. Such strategies are unfortunately not explicitly built into the backbone of the present-day language model. This criticism has been floating around as an informal viewpoint (LeCun, 2024; Bubeck et al., 2023). Our paper is aimed at crystallizing this intuitive criticism of next-token prediction, and developing the core arguments of this debate.
@@
Although this argument is compelling, a pessimist would worry, realistically, even with minor imperfections in the next-token predictor, the accuracy may break down spectacularly for long sequences (Kääriäinen, 2006; Ross & Bagnell, 2010; LeCun, 2024; Dziri et al., 2023). Say, even if every next-token error is as little as $0.01$, the probability of encountering an erroneous token exponentially compounds along the way, and by the end of $200$ tokens, blows up to $0.86$.
@@
This is a simple and powerful observation. Yet, this does not completely capture the intuition that next-token predictors may be poor planners. Crucially, this argument does not carefully distinguish between the two types of next-token prediction: inference-time autoregression (where the model consumes its own previous outputs as inputs), and training-time teacher-forcing (Williams & Zipser, 1989) (where the model is taught to predict token-by-token consuming all previous ground truth tokens as inputs). Framed this way, the compounding of errors only pinpoints a superficial failure to execute a plan during inference. It leaves open the possibility that we may have still learned a near-perfect next-token predictor; perhaps, with an appropriate post-hoc wrapper that verifies and backtracks, we can elicit the right plan without compounding errors.
@@
Drawing this distinction allows us to articulate a much more concerning possibility: is it safe to assume that next-token based learning (teacher-forcing) always learns an accurate next-token predictor? We identify this is not always the case. Consider a task where we expect the model to witness a problem statement $\bm{p}=(p_{1},p_{2}\ldots,)$ and produce the ground truth response tokens $(r_{1},r_{2},\ldots)$. Teacher-forcing trains the model to produce each token $r_{i}$ by not only providing the problem statement $\bm{p}$ but also by revealing part of the ground truth $r_{1},\ldots r_{i-1}$. Depending on the task, we first argue that this can induce shortcuts that use the revealed prefix of the ground truth answer to spuriously fit future answer tokens. We call this the Clever Hans cheat.  [^1] Clever Hans (Pfungst & Rahn, 1911) was a famous show horse that could solve simple arithmetic tasks by repeatedly tapping with his hoof until he reached the correct count. It turns out however, that Clever Hans did not really solve the problem, but merely stopped tapping upon detecting certain (involuntary) facial cues from his coach. Clever Hans’ answers were wrong when the coach was absent. Next, while the later tokens ($r_{i}$ for large $i$) become easy to fit by the Clever Hans cheat, in contrast, the earlier answer tokens (say, $r_{0},r_{1}$ etc.,) become harder to learn. This is because they no longer come with any supervision about the full answer — part of the supervision is lost to the Clever Hans cheat. We argue that these two flaws would together arise in “lookahead tasks”: tasks that require implicitly planning a later token in advance of an earlier token. In such tasks, teacher-forcing would result in a highly inaccurate next-token predictor that would fail to generalize to unseen problems $\bm{p}$, even those sampled in-distribution.
@@
Empirically, we demonstrate that the above mechanism leads to complete in-distribution failure in a path-finding setup on a graph, that we propose as a minimal lookahead task. We design our setup in a way that it is demonstrably straightforward to solve that the failure of any model is remarkable. Yet, we observe failure for both the Transformer (Vaswani et al., 2017) and the Mamba architecture, a structured state space model (Gu & Dao, 2023). We also find that a form of teacherless training that predicts multiple future tokens (Monea et al., 2023) is (in some settings) able to circumvent this failure. Thus, we pinpoint a precise and easy-to-learn scenario where, rather than properties that are criticized in existing literature — like convolution or recurrence or autoregressive inference (see 6), — it is next-token prediction during training that is at fault.
@@
    We identify that the next-token prediction debate must not conflate autoregressive inference with teacher-forcing. Both lead to vastly different failures (§3A).
@@ 2 The Two Modes of Next-Token Prediction
(Inference-time next-token prediction via autoregression) Autoregressive inference is a form of inference-time next-token prediction in that to generate a response $\hat{\bm{r}}$, we iterate over $i=1,\ldots,L_{\texttt{resp}}$, to sample the next token $\hat{r}_{i}$ with the distribution given by ${\texttt{LM}}_{\theta}(\hat{r}_{i}\;;\bm{p},\hat{\bm{r}}_{<i})$. We denote this as $\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}{\texttt{LM}}_{\theta}(\cdot\;;\bm{p})$.
@@ 3 Failure due to auto-regressive inference
(Every sequence distribution can be represented by a next-token predictor) By the chain rule of probability we have ${\mathbb{P}}_{\mathcal{D}}(\bm{r}\;|\;\bm{p})=\prod_{i=1}^{L_{\texttt{resp}}}{\mathbb{P}}_{\mathcal{D}}(r_{i}\;|\;\bm{p},\bm{r}_{<i})$. Therefore, define a next-token predictor LM such that for every valid value of $i,\bm{p}$, and $\bm{r}$, we have $\texttt{LM}(\hat{r}_{i}=r_{i}\;;\bm{p},\bm{r}_{<i})\coloneq{\mathbb{P}}_{\mathcal{D}}(r_{i}|\bm{p},\bm{r}{<i})$. Then, sampling $\bm{r}\sim\mathcal{D}|\bm{p}$, is equivalent to autoregressively sampling $\bm{r}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}(\cdot\;;\bm{p})$.
@@
The snowballing errors criticism: A skeptic would however raise the following concern. Regardless of the abundance of computational resources, realistic models are typically not perfect next-token predictors. There may always be a slight chance of error in each step, and once an error is committed, there is no explicit backtracking mechanism to rescue the model. The argument then goes that, the probability of errors in each token, however miniscule, would exponentially snowball along the way. By the end of a long sequence of tokens, the accuracy (of having produced an error-free response) becomes trivial. This has been formalized in various contexts, from that of autoregressive models LeCun (2024), to that of the limits of Transformers in compositional tasks Dziri et al. (2023), and in a different form, in much earlier work in imitation learning and structured prediction Kääriäinen (2006); Ross & Bagnell (2010) (see §6). We present a minimal formalization of this below:
@@
Then, for $\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot\;;\bm{p})$ the probability that the generated response exactly matches the ground truth $\bm{r}$ obeys
@@
We argue that the snowball failure mode only indicates how an autoregressive model can fail to execute a plan during inference-time. It does not preclude the possibility that the model may have learned a good plan that it simply fails to execute during inference. Concretely, it may still be possible that, at each step, the model has high accuracy of predicting a next token that is consistent with a good plan (as assumed in Eq 3). Depending on the setting, one can potentially exploit this accuracy to elicit a good plan during inference. For instance, one may be able to use a post-hoc wrapper that first verifies whether an error has taken place, then backtracks and re-executes a different action. One may even simulate backtracking using more elaborate techniques such as chain/tree/graph of thought (Wei et al., 2022; Yao et al., 2023a; Besta et al., 2023; Yao et al., 2023b), or using the model to give itself feedback (Madaan et al., 2023; Huang et al., 2022; Shinn et al., 2023) to elicit the plan that the model has learned.
@@ 4.2 Outline of failure mechanism
While we will use the path-star example as a running example, we make our claim more generally for problems that require lookahead (such as story-writing, as we will discuss later). In such classes of problems, we claim that teacher-forcing prevents learning the true mechanisms, causing failure. Intuitively, in teacher-forcing, we decompose the learning of $\bm{p}\to\bm{r}$ into multiple problems, one for each token $r_{i}$. Specifically, we make the model learn a mapping from the input $(\bm{p},\bm{r}_{<i})$ — not just $\bm{p}$ — to the output $r_{i}$. The additional information $\bm{r}_{<i}$ in the input, we argue, is problematic and destroys the core challenge in what the model has to learn. Specifically, our argument puts forth two debilitating mechanisms that would together emerge under teacher-forcing (explained over the next two subsections). While, we will empirically verify these mechanisms for path-star graphs in §5, we also provide a discussion of how our ideas apply to a text-based scenario at the end of this section.
@@ 4.3 The Clever Hans Cheat
To illustrate this in our path-star example, without loss of generality, consider a ground truth path that is of the form $\bm{r}={v}_{\tt{start}},v_{1},v_{2},\ldots,{v}_{\tt{goal}}$. With a slight abuse of the indexing notation, let $\bm{r}_{<i}={v}_{\tt{start}},v_{1},\ldots,v_{i-1}$ be the prefix of length $i$ (so we index from 00 instead of $1$). Observe that nodes from $v_{2}$ onwards, until before ${v}_{\tt{goal}}$, have precisely one edge going “away” from ${v}_{\tt{start}}$. Thus, consider when the model is given as input, $(\bm{p},\bm{r}_{<i})$ where $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$, to fit the target $v_{i}$. The model first merely needs to scan the adjacency list $\texttt{adj}(G)$ within $\bm{p}$ for the one edge containing $v_{i-1}$ in the first position. Then, the model only has to predict the other node on that edge as $v_{i}$. Note though, this cheat cannot work on fitting the target $v_{1}$ given the input $\bm{r}_{<1}={v}_{\tt{start}}$ since ${v}_{\tt{start}}$ has many outward edges — we will address this node in the next section. We illustrate this difference between $\bm{r}_{1}$ and the remaining tokens as the “easy” vs. “difficult” tokens in Fig. 2.
@@
We emphasize two key aspects of this cheating behavior. First, these shortcuts are unlike previously-identified shortcuts (see §6) that map from the original input prefixes $\bm{p}$ to the ground truth $\bm{r}$. The behavior we identify is unique to the mapping from the teacher-forced prefix $\bm{p},\bm{r}_{<i}$ to $r_{i}$. We name this behavior as Clever Hans cheating. Another notable point is that this does not come from a dearth of samples: even if we had infinite training data at our disposal, the model can still fit the easy tokens of all that data by Clever Hans cheating.
@@ 4.4 The Indecipherable Token
Perhaps, not all is lost. While the later tokens may be fit using the Clever Hans cheat, we may still have some of the earlier tokens (for small $i$), for which such cheats may be unavailable. The supervision from these tokens may eventually coerce the model into learning the true solution. For example, in the path-star task, the model still needs to learn to predict the first node $v_{1}$, where it is not possible to fit the training data by the Clever Hans cheat. If not memorize this token on the training data, the most general way to fit this token is by actually solving the underlying task.
@@
However, we argue that it is significantly harder for the model to learn the correct solution now. Consider the point in training when the Clever Hans cheat is perfected. At this point, the model is deprived of information about much of the full solution which was once present as supervisory targets. The model is simply left with the task of mapping the input $\bm{p}$ to an incomplete solution (e.g., the first vertex $v_{1}$ in the path-star graph). Recovering the plan in this scenario must first of all be relatively harder from a statistical point of view due to the incomplete supervision. But more importantly, learning this task may become computationally harder, or simply, intractable (Wies et al., 2023).
@@
We provide an informal intuition of intractability for the path-star problem, but this intuition should extend to more general problems as well. Intuitively, our learner has to find an end-to-end algorithm that composes multiple subroutines over one another. For instance, recall that the straightfoward solution consists of $l$ steps: start from the current vertex as ${v}_{\tt{goal}}$, and find the preceding vertex in the graph in each subsequent step. Each vertex in this path can be thought of as “intermediate supervision” to learn a corresponding “find-the-adjacent-vertex” subroutine from a space of candidate subroutines. [^2] As an illustration of what these candidate subroutines could be, imagine that the model can implement an induction head (Olsson et al., 2022) $\texttt{Ind}_{k}(\bm{p},v)$ that finds $v$ in the adjacency list of $\bm{p}$, and outputs the token that precedes it by $k$ positions. Then the candidate space could be parameterized by $k$ as $\{\texttt{Ind}_{k}(\bm{p},v)|k=1,2,\ldots,\}$. For our specific tokenization, the correct subroutine at each of the $l$ steps is the induction head for which $k=2$. Even if we conservatively assume that there is only a constant-sized space of $C$ candidate subroutines, the end-to-end search space is an exponential space of $C^{l}$ algorithms composing $l$ subroutines.
@@
Now, after the Clever Hans cheat is in effect, the only supervision for this search is a single-token loss of the form $-\log{\texttt{LM}}_{\theta}\left(\hat{r}_{1}={r}_{1};\bm{p}\right)$. However, this loss is an “all-or-nothing” loss. Crucially, by the discrete nature of the task, even if the learner gets one subroutine incorrect, the final answer $\hat{r}_{1}$ would likely be incorrect on all inputs. For instance, imagine that the first subroutine is incorrectly learned and its output takes us to an arbitrary location on the graph. Then, even if all subsequent subroutines were correctly learned (i.e., they are “find-the-adjacent-vertex” subroutines), the final output would be arbitrary. Thus, we have $\hat{r}_{1}={r}_{1}$ precisely for the algorithm where all $l$ subroutines are correct, and $\hat{r}_{1}\neq{r}_{1}$ for any other choice of the algorithm. For such an all-or-nothing loss surface, the end-to-end learner must necessarily brute-force search through the exponential space of algorithms.  [^3] The only condition under which the learner may identify the right algorithm efficiently is when the model has accidentally seen similar problems during pretraining and thus has a useful prior. For example, the prior may be that it assigns high probability to the correct subroutine(s). Or, in the specific path-star example, the model may assign high probability to all $l$ subroutines being identical. In such a case, one can construct slight variations of the tasks that defy this prior, to demonstrate intractability.
We encapsulate the overall claim more generally via the following conditions:
@@
Consider learning a task that satisfies the following conditions:
  1. 1.
    It requires composing $l$ discrete-output subroutines over one another.
  2. 2.
    The $k$ leading response tokens are sensitive in that even if one subroutine is altered, the first $k$ tokens are each completely altered.
  3. 3.
    The subroutine search space consists of at least $C$ candidate subroutines.
Then learning the task with only supervision from the first $k$ ground truth tokens requires exponential time of $\Omega(C^{l})$.
Indeed, literature on chain-of-thought has identified many instances of similar “multi-hop reasoning tasks” that cannot be solved end-to-end — both empirically (Wei et al., 2022) and theoretically (Wies et al., 2023) — unless there is “complete supervision”. We elaborate on this in §6.
In § 5, we will verify experiments to demonstrate that our models indeed fail to learn the Indecipherable Token as a result of Clever Hans cheating, and that conversely, they succeed in conditions where the Clever Hans cheat is prevented.

### v2

@@ Abstract
Can a mere next-token predictor faithfully model human intelligence? We crystallize this emerging concern and correct popular misconceptions surrounding it, and advocate a simple multi-token objective. As a starting point, we argue that the two often-conflated phases of next-token prediction — autoregressive inference and teacher-forced training — must be treated distinctly. The popular criticism that errors can compound during autoregressive inference, crucially assumes that teacher-forcing has learned an accurate next-token predictor. This assumption sidesteps a more deep-rooted problem we expose: in certain classes of tasks, teacher-forcing can simply fail to learn an accurate next-token predictor in the first place. We describe a general mechanism of how teacher-forcing can fail, and design a minimal planning task where both the Transformer and the Mamba architecture empirically fail in that manner — remarkably, despite the task being straightforward to learn. Finally, we provide preliminary evidence that this failure can be resolved using a simple modification that predicts multiple tokens in advance. We hope this finding can ground future debates and inspire explorations beyond the next-token prediction paradigm. We make our code available under https://github.com/gregorbachmann/Next-Token-Failures
Machine Learning, ICML
@@ 1 Introduction
Long after its inception in the seminal work of Shannon (1948, 1951), next-token prediction has made its way into the core of the modern language model. But despite its long list of achievements, there is a small but growing belief that a next-token predicting model is merely an impressive improv artist that cannot truly model human thought. Humans, when navigating the world, meticulously imagine, curate and backtrack plans in their heads before executing them. Such strategies are unfortunately not explicitly built into the backbone of the present-day language model. This criticism has been floating around as an informal viewpoint (LeCun, 2024; Bubeck et al., 2023). Our paper is aimed at crystallizing this intuitive criticism, clarifying popular misconceptions, and developing new core arguments for the next-token prediction debate.
@@
Although this argument is compelling, a pessimist would worry, realistically, even with minor imperfections in the next-token predictor, the accuracy may fall spectacularly for long sequences (Kääriäinen, 2006; Ross & Bagnell, 2010; LeCun, 2024; Dziri et al., 2024). Say, even if every next-token error is as little as $0.01$, the probability of encountering an erroneous token exponentially compounds along the way, and by the end of $200$ tokens, blows up to $0.86$.
@@
This is a simple and powerful observation. Yet, we explain why this does not completely capture the intuition that next-token predictors may be poor planners. Crucially, this argument does not carefully distinguish between the two types of next-token prediction: inference-time autoregression (where the model consumes its own previous outputs as inputs), and training-time teacher-forcing (Williams & Zipser, 1989) (where the model is taught to predict token-by-token consuming all previous ground truth tokens as inputs). Framed this way, the compounding of errors only pinpoints a superficial failure to execute a plan during inference. It leaves open the possibility that we may have still learned a near-perfect next-token predictor; perhaps, with an appropriate post-hoc wrapper that verifies and backtracks, we can elicit the right plan without compounding errors.
@@
Drawing this distinction allows us to articulate a much more concerning possibility: is it safe to assume that next-token based learning (teacher-forcing) always learns an accurate next-token predictor? We identify this is not always the case. Consider a task where we expect the model to witness a problem statement $\bm{p}=(p_{1},p_{2}\ldots,)$ and produce the ground truth response tokens $(r_{1},r_{2},\ldots)$. Teacher-forcing trains the model to produce each token $r_{i}$ by not only providing the problem statement $\bm{p}$ but also by revealing part of the ground truth $r_{1},\ldots r_{i-1}$. Depending on the task, we first argue that this can induce shortcuts that use the revealed prefix of the ground truth answer to spuriously fit future answer tokens. We call this the Clever Hans cheat.  [^1] Clever Hans (Pfungst & Rahn, 1911) was a famous show horse that could solve simple arithmetic tasks by repeatedly tapping with his hoof until he reached the correct count. It turns out however, Clever Hans did not really solve the problem, but merely stopped tapping upon detecting certain (involuntary) cues from his coach. Clever Hans’ answers were wrong when the coach was absent. Next, while the later tokens ($r_{i}$ for large $i$) become easy to fit by the Clever Hans cheat, in contrast, the earlier answer tokens (say, $r_{0},r_{1}$ etc.,) become harder to learn. This is because they no longer come with any supervision about the full answer — part of the supervision is lost to the Clever Hans cheat. We argue that these two flaws would together arise in “lookahead tasks”: tasks that require implicitly planning a later token in advance of an earlier token. In such tasks, teacher-forcing would result in a highly inaccurate next-token predictor that would struggle to generalize to unseen problems $\bm{p}$, even those sampled in-distribution.
@@
Empirically, we demonstrate that the above mechanism leads to complete in-distribution failure in a path-finding setup on a graph, that we propose as a minimal lookahead task. We design our setup in a way that it is demonstrably straightforward to solve that the failure of any model is remarkable. Yet, we observe failure for both the Transformer (Vaswani et al., 2017) and the Mamba architecture, a structured state space model (Gu & Dao, 2023). We also find that a form of teacherless training that predicts multiple future tokens (Monea et al., 2023) is (in some settings) able to circumvent this failure. Thus, we pinpoint a precise and easy-to-learn scenario where, rather than properties that are criticized in existing literature — like convolution or recurrence or autoregressive inference (see §6), — it is next-token prediction during training that is at fault.
@@
    We identify that the next-token prediction debate must not conflate autoregressive inference with teacher-forcing. Both lead to vastly different failures (§3B).
@@ 2 The Two Modes of Next-Token Prediction
(Inference-time next-token prediction via autoregression) Autoregressive inference is a form of inference-time next-token prediction in that to generate a response $\hat{\bm{r}}$, we iterate over $i=1,\ldots,L_{\texttt{resp}}$, to sample the next token $\hat{r}_{i}$ with the distribution given by ${\texttt{LM}}_{\theta}(\hat{r}_{i}\;;\bm{p},\hat{\bm{r}}_{<i})$. We denote this as $\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}{\texttt{LM}}_{\theta}(\cdot\;;\bm{p})$.
@@ 3 Failure due to Auto-Regressive Inference
(Every sequence distribution can be represented by a next-token predictor) By the chain rule of probability we have ${\mathbb{P}}_{\mathcal{D}}(\bm{r}\;|\;\bm{p})=\prod_{i=1}^{L_{\texttt{resp}}}{\mathbb{P}}_{\mathcal{D}}(r_{i}\;|\;\bm{p},\bm{r}_{<i})$. Therefore, define a next-token predictor LM such that for every valid value of $i,\bm{p}$, and $\bm{r}$, we have $\texttt{LM}(\hat{r}_{i}=r_{i}\;;\bm{p},\bm{r}_{<i})\coloneq{\mathbb{P}}_{\mathcal{D}}(r_{i}|\bm{p},\bm{r}{<i})$. Then, sampling $\bm{r}\sim\mathcal{D}|\bm{p}$, is equivalent to autoregressively sampling $\bm{r}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}(\cdot\;;\bm{p})$.
@@
The snowballing errors criticism: A skeptic would however raise the following concern. Regardless of the abundance of computational resources, realistic models may still predict the next token with a slight probability of error in each step; these error probabilities may then exponentially accumulate over time. This has been formalized in various contexts, from that of autoregressive models LeCun (2024), to that of the limits of Transformers in compositional tasks Dziri et al. (2024), and in a different form, in much earlier work in imitation learning and structured prediction Kääriäinen (2006); Ross & Bagnell (2010) (see §6). We present a minimal formalization of this below:
@@
Then, for $\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot\;;\bm{p})$ the probability that the generated response exactly matches the ground truth $\bm{r}$ obeys
@@
We argue that the snowball failure mode only indicates how an autoregressive model can fail to execute a plan during inference-time. It does not preclude the possibility that the model may have learned a good plan that it simply fails to execute during inference. Concretely, it may still be possible that, at each step, the model has high accuracy of predicting a next token that is consistent with a good plan (as assumed in Eq 3). Depending on the setting, one can potentially exploit this accuracy to elicit a good plan during inference. For instance, one may be able to use a post-hoc wrapper that verifies whether an error has taken place, then backtracks and executes a different action. One may even simulate backtracking using more elaborate techniques such as tree-of-thought (Wei et al., 2022; Yao et al., 2023a; Besta et al., 2024; Yao et al., 2023b), or using the model to give itself feedback (Madaan et al., 2024; Huang et al., 2022; Shinn et al., 2023) to elicit the plan that the model has learned.
@@ 4.2 Outline of Failure Mechanism
While we will use the path-star example as a running example, we make our claim more generally for problems that require lookahead (such as story-writing, as we will discuss later). In such problems, we claim that teacher-forcing prevents learning the true mechanisms, causing failure. Intuitively, in teacher-forcing, we decompose the learning of $\bm{p}\to\bm{r}$ into multiple problems, one for each token $r_{i}$. Specifically, we make the model learn a mapping from the input $(\bm{p},\bm{r}_{<i})$ — not just $\bm{p}$ — to the output $r_{i}$. The additional information $\bm{r}_{<i}$ in the input, we argue, is problematic and destroys the core challenge in what the model has to learn. Specifically, our argument puts forth two debilitating mechanisms that would together emerge under teacher-forcing (explained over the next two subsections). While, we will empirically verify these mechanisms for path-star graphs in §5, we also provide a discussion of how our ideas apply to a text-based scenario at the end of this section.
@@ 4.3 The Clever Hans Cheat
To illustrate this in our path-star example, without loss of generality, consider a ground truth path that is of the form $\bm{r}={v}_{\tt{start}},v_{1},v_{2},\ldots,{v}_{\tt{goal}}$. With a slight abuse of the indexing notation, let $\bm{r}_{<i}={v}_{\tt{start}},v_{1},\ldots,v_{i-1}$ be the prefix of length $i$ (so we index from 00 instead of $1$). Observe that nodes from $v_{2}$ onwards, until before ${v}_{\tt{goal}}$, have precisely one edge going “away” from ${v}_{\tt{start}}$. Thus, consider when the model is given as input, $(\bm{p},\bm{r}_{<i})$ where $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$, to fit the target $v_{i}$. The model first merely needs to scan the adjacency list $\texttt{adj}(G)$ within $\bm{p}$ for the one edge containing $v_{i-1}$ in the first position. Then, the model only has to predict the other node on that edge as $v_{i}$. Note though, this cheat cannot work on fitting the target $v_{1}$ given the input $\bm{r}_{<1}={v}_{\tt{start}}$ since ${v}_{\tt{start}}$ has many outward edges — we will address this node in the next section. We illustrate this difference between $v_{1}$ and the remaining tokens as the “easy” vs. “difficult” tokens in Fig. 2.
@@
We emphasize two key aspects of this cheating behavior. First, these shortcuts are unlike well-known shortcuts (see Remark 5) that map from the original input prefixes $\bm{p}$ to the ground truth $\bm{r}$. What we identify is unique to the mapping from the teacher-forced prefix $(\bm{p},\bm{r}_{<i})$ to $r_{i}$. We name this Clever Hans cheating. Another notable point is that this does not come from a dearth of samples: even if we had infinite training data at our disposal, the model can still fit the easy tokens of all that data by Clever Hans cheating.
@@ 4.4 The Indecipherable Token
Perhaps, not all is lost. While the later tokens may be fit using the Clever Hans cheat, we may still have some of the earlier tokens (for small $i$), for which such cheats may be unavailable. The supervision from these tokens may eventually coerce the model into learning the true solution. For example, in the path-star task, the model still needs to learn to predict the first node $v_{1}$, where it is not possible to fit the training data by the Clever Hans cheat. If not memorize this token on the data, the most general way to fit this token is by actually solving the underlying task.
@@
However, we argue that it is significantly harder for the model to learn the correct solution now. Consider the moment in training when the Clever Hans cheat is perfected. At this point, the model is deprived of information about much of the full solution which was once present as supervisory targets. The model is simply left with the task of mapping the input $\bm{p}$ to an incomplete solution (e.g., the first vertex $v_{1}$ in the path-star graph). Recovering the plan in this scenario must first of all be relatively harder from a statistical point of view due to the incomplete supervision. But more importantly, learning this task may become computationally harder, or even intractable under certain assumptions. We provide an informal intuition using the path-star problem below, but this intuition should extend to more general problems as well — indeed, Wies et al. (2023) and other literature on chain-of-thought echo similar negative results about learning from limited supervision (see §H).
@@
Intuitively, our learner has to find an end-to-end algorithm that composes multiple subroutines. For instance, the straightfoward solution consists of $l$ steps: start from the current vertex as ${v}_{\tt{goal}}$, and find the preceding vertex in the graph in each subsequent step. Each vertex in this path can be thought of as “intermediate supervision” to learn a corresponding “find-the-adjacent-vertex” subroutine from a space of candidate subroutines. [^2] As an illustration of what these candidate subroutines could be, imagine that the model can implement an induction head (Olsson et al., 2022) $\texttt{Ind}_{k}(\bm{p},v)$ that finds $v$ in the adjacency list of $\bm{p}$, and outputs the token that precedes it by $k$ positions. Then the candidate space could be parameterized by $k$ as $\{\texttt{Ind}_{k}(\bm{p},v)|k=1,2,\ldots,\}$. For our specific tokenization, the correct subroutine at each of the $l$ steps is the induction head for which $k=2$. Even if we conservatively assume that there is only a constant-sized space of candidate subroutines $\mathcal{C}$, the end-to-end search space is an exponential space of $|\mathcal{C}|^{l}$ algorithms composing $l$ subroutines.
@@
Now, after the Clever Hans cheat is in effect, the only supervision for this search is the single-token loss, $-\log{\texttt{LM}}_{\theta}\left(\hat{r}_{1}={r}_{1};\bm{p}\right)$. However, this loss is an “all-or-nothing” loss. Crucially, by the discrete nature of the task, even if one subroutine is incorrect, the final answer $\hat{r}_{1}$ would likely be incorrect on all inputs. For instance, imagine that the first subroutine is incorrect and its output takes us to an arbitrary location on the graph. Then, even if all subsequent subroutines were correct (i.e., they are “find-the-adjacent-vertex” subroutines), the final output would be arbitrary. Thus, we have $\hat{r}_{1}={r}_{1}$ precisely for the algorithm where all $l$ subroutines are correct, and $\hat{r}_{1}\neq{r}_{1}$ for any other choice of the algorithm. For such an all-or-nothing loss surface, the end-to-end learner must necessarily brute-force search the exponential space of algorithms. We encapsulate the overall claim more generally below:
@@
Let $\mathcal{C}$ be a set of discrete-output candidate subroutines. Consider learning a task such that (i) it requires composing some $l$ subroutines from $\mathcal{C}$, (ii) the $k$ leading response tokens are sensitive in that even if one subroutine is altered, the first $k$ tokens are each completely altered. Then learning the task with only supervision from the first $k$ ground truth tokens requires exponential time of $\Omega(|\mathcal{C}|^{l})$, given the assumptions listed in Remark 1.
In §5, we will verify experiments to demonstrate that our models indeed fail to learn the Indecipherable Token as a result of Clever Hans cheating, and that conversely, they succeed whenever the Clever Hans cheat is prevented.